# Improving Parallel Program Performance Through DSL-Driven Code Generation with LLM Optimizers

### v1

@@ Abstract
Mapping computations to processors and assigning data to memory are critical for maximizing performance in parallel programming. These mapping decisions are managed through the development of specialized low-level system code, called mappers, crafted by performance engineers. Each mapper is tailored to a specific application and optimized for the underlying machine architecture, a process that requires days of refinement and tuning from an expert. Despite advances in system research, automating mapper generation remains a challenge due to the complexity of making millions of decisions to find the optimal solution and generate the solution as code. We introduce an approach that leverages recent advances in LLM-based optimizers for mapper design. In under ten minutes, our method automatically discovers mappers that surpass human expert designs in scientific applications by up to 1.34×1.34$\times$1.34 × speedup. For parallel matrix multiplication algorithms, our mapper achieves up to 1.31×1.31$\times$1.31 × of the expert-designed solution. To achieve this, we simplify the complexity of low-level code generation by introducing a domain-specific language (DSL) that abstracts the low-level system programming details and defines a structured search space for LLMs to explore. To maximize the application performance, we use an LLM optimizer to improve an agentic system that generates the mapper code. As a result, this approach significantly reduces the workload for performance engineers while achieving substantial performance gains across diverse applications. Finally, our results demonstrate the effectiveness of LLM-based optimization in system design and suggest its potential for addressing other complex system challenges.
@@ 1 Introduction
Task-based programming (Slaughter et al., 2015; Bauer et al., 2012; Augonnet et al., 2009; Chamberlain et al., 2007; Kaiser et al., 2014; Heller et al., 2017; Chandra et al., 2001; Duran et al., 2011; Moritz et al., 2018; Barham et al., 2022) has recently emerged as a prominent paradigm in parallel programming. The core idea is to decompose computations into self-contained functions called tasks that do not communicate with other tasks except through their arguments (typically tensors or multi-dimensional arrays). A performance-critical aspect of executing task-based applications is mapping, which involves assigning tasks to processors, mapping data to physical memories, and managing other low-level hardware resources. A concrete mapping policy, implemented as code, is referred to as a mapper. Mappers focus solely on improving performance without altering the correctness of the application’s output. The performance impact of a good versus a poor mapper can be substantial, often resulting in differences of an order of magnitude (Abdel-Gawad et al., 2014; Galvez et al., 2017).
Currently, writing mappers is a manual, labor-intensive, and time-consuming process that requires deep knowledge of the application, the machine, and the low-level system’s C++ mapping APIs. Even for experienced engineers, developing a high-performance mapper can take several days, requiring multiple rounds of refinement (SFX Teixeira et al., 2023).
In this paper, we focus on using LLMs to automatically generate mappers. There are two main challenges in using LLMs for mapper code generation. First, given a specific mapping policy described in natural language, LLMs struggle to generate the corresponding mapper code, which consists of a few hundred lines of low-level C++ code. This code generation task is significantly different from function-level code generation, which is more isolated and modular, where LLMs have demonstrated strong performance (Chen et al., 2021; Austin et al., 2021). In our case, generating C++ mappers requires a deep understanding of low-level systems and broader contextual dependencies in large-scale system code, which is beyond the capabilities of current LLMs (Jimenez et al., 2023; Du et al., 2023; Wang et al., 2024). The second challenge lies in identifying the optimal mapping strategy to maximize application performance. Many design problems in systems are naturally constrained optimization problems. Generating mappers that are syntactically correct is insufficient.
To address the first challenge, we design a Domain-Specific Language (DSL) for mapper generation. The DSL provides a high-level interface that encapsulates all the critical decisions a mapper has to make without interacting with the intricate low-level C++ APIs. The DSL defines a search space in its abstractions, which breaks a monolithic code generation down to generating multiple independent, modular components, which is much easier for code generation due to less contextual dependence between components.
For the second challenge, oirst, structuring the search space through the DSL’s decomposition and modularization of the mapper, and second, applying reinforcement learnformulating the search problem as an optimization problem where the objective is to maximize the performance of the parallel program. Similar approaches have achieved significant results in discovering faster algorithms for matrix multiplication (AlphaTensor) (Fawzi et al., 2022) and generating mathematical proofs (AlphaGeometry) (Trinh et al., 2024), both require searching through a discrete space for an optimal solution.
Recently, a great deal of work has demonstrated that instruction-tuned LLMs have an inherent ability to “optimize” for an objective, such as minimizing a function (Nie et al., 2024; Cheng et al., 2024), prompt optimization (Yang et al., ; Khattab et al., 2023), and molecule design optimization (Yuksekgonul et al., 2024). We propose a fully automated agent-based system to generate mappers and optimize for parallel application performance, illustrated in Figure 1. Our solution allows LLMs to refine their mapping strategies efficiently, reducing the time required to develop high-performance mappers from days to just a few minutes.
Figure 1: We illustrate an agent-based optimization for high-performance mapper code generation, where the mapper is written in a Domain-Specific Language (DSL) and iteratively refined by an agent through performance feedback. Initially, all tasks are mapped to the CPU and system memory. The LLM then improves by mapping computation-intensive tasks to the GPU and GPU’s FrameBuffer memory to minimize task execution time. Finally, tasks that require significant data exchange between the CPU and GPU are mapped to the ZeroCopy memory, enabling shared access and minimizing data movement overhead.
Our experiments show that mappers generated by LLMs not only match but can surpass expert-written mappers, achieving up to 1.34×1.34$\times$1.34 × speedup across 9 benchmarks. Notably, the search process enhances the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×. We also demonstrate, through ablation studies, the critical role of high-quality feedback in guiding LLM optimizers to discover more efficient mapping strategies. Our contributions are as follows:
@@
    Development of a Domain-Specific Language (DSL) to Simplify Mapper Generation: We design a DSL to address the complexities of direct code generation in low-level systems. This DSL provides a higher-level abstraction, enabling LLMs to generate mapping code more effectively by encapsulating the complexities of low-level C++ APIs.
@@
    Formulation of Mapper Generation as a Discrete Optimization Problem: We formulate the task of mapper code generation as a discrete optimization problem with the objective of maximizing application performance. Our DSL defines a more structured and constrained search space for LLM to explore.
@@
    Experimental Validation and Performance Comparison: Our DSL significantly improves the success rate of mapper code generation for a given mapping strategy from 0% in C++ to 80% in our DSL. Leveraging high-quality feedback, our agent-based solution discovers mappers that achieve up to 1.34×1.34$\times$1.34 × speedup compared with expert-designed mappers, enhancing the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×.
@@ 2 Related Work
Mapping in Parallel Programming Many parallel programming systems allow users to make their own mapping decisions, such as Legion (Bauer et al., 2012), StarPU (Augonnet et al., 2009; 2010), Chapel (Chamberlain et al., 2007), HPX (Kaiser et al., 2014; Heller et al., 2017), Sequoia (Fatahalian et al., 2006), Ray (Moritz et al., 2018), and Pathways (Barham et al., 2022). Several techniques have been proposed to automate mapping, including machine learning models (O’Boyle et al., 2013; Wang & O’Boyle, 2009), static analysis (Poesia et al., 2017; Ren et al., 2008), and auto-tuning (SFX Teixeira et al., 2023). Unlike previous work, we use an agent-based approach with LLMs, exploring a larger search space for mappers than traditional methods.
Automated Code Generation With the rise of LLMs, various models (Chen et al., 2021; Nijkamp et al., 2022; Li et al., 2023; Ouyang et al., 2022; Wei et al., 2023) have been trained on code, enabling them to assist with programming tasks and automate software development (Shypula et al., 2023; Yang et al., 2024b; Jain et al., 2024). While LLMs have shown success in generating isolated function-level Python code (Chen et al., 2021; Austin et al., 2021), they face challenges when generating system-level C++ code in large repositories due to complex APIs and broader contextual dependencies (Du et al., 2023; Wang et al., 2024). Thus, adopting LLMs to generate code in real-world software remains difficult (Jimenez et al., 2023). Our work addresses this by developing a domain-specific language (DSL) that provides a high-level abstraction to help LLMs generate mapper code more effectively.
Automated System Design Optimization The use of machine learning for optimizing system design has garnered significant attention in recent years. Learned cost models have proven effective in optimizing deep learning operators on GPUs (Zheng et al., 2020a; b; 2022b). Methods in reinforcement learning and multi-armed bandit have been applied to tasks such as chip floorplanning (Mirhoseini et al., 2021) and program autotuning (Ansel et al., 2014), automatic vectorization of programs (Haj-Ali et al., 2020a), and optimizing compiler phase ordering (Haj-Ali et al., 2020b). Previous work primarily focuses on hyperparameter configurations, while our work aims to improve system performance by directly generating mapper code.
Optimization with LLMs
Recently, LLMs have been increasingly applied to optimization problems traditionally found in numerical domains (Boyd & Vandenberghe, 2004). AhmadiTeshnizi et al. (2023) addressed structured optimization challenges, such as mixed-integer linear programming (LP), while Yang et al. (2024a) applied LLMs to unstructured problems, treating them as black-box optimizations. (Nie et al., 2024) highlighted the role of feedback in efficiently finding global minima for black-box functions. Cheng et al. (2024) formally defined optimization problems solvable by LLMs and proposed an optimizer based on execution graphs. More recently, Yuksekgonul et al. (2024) introduced an LLM optimizer inspired by gradient descent. These advancements demonstrate the potential of LLMs to optimize beyond task completion, improving performance metrics across various domains.
3 Mapper Generation Task
Task Definition
The problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012). Mappers determine how tasks are assigned to processors and how data is placed in memory to optimize performance. A well-designed mapper can achieve up to 10$\times$ speedup compared to random mapping strategies. However, mappers are highly context-specific and must be carefully tailored to an application’s input and the machine’s architecture. Finding an optimal mapper is akin to solving a combinatorial optimization problem—a process that is time-consuming, labor-intensive, and traditionally performed by experienced performance engineers. The search space for discovering the best mapping strategy is vast, growing to $2^{14}$ even for the simplest scientific applications (as shown in prior work with a smaller search space than ours (SFX Teixeira et al., 2023)), making it challenging to efficiently explore all possible solutions. Moreover, even with a clear mapping strategy, writing the corresponding mapper requires experts to produce hundreds of lines of low-level C++ code, a task that can take several hours. As a result, the entire performance tuning process can take several days to complete.
Figure 2: Mappers decide the placement of each task in the task graph to processors, the placement of data to memory, and how the iteration space of data is partitioned and mapped to different processors.
Mapping Decisions
@@ 3 Mapper Generation Task
Now we elaborate on the decisions that a mapper has to make. The first critical decision is the processor selection for each task, determining whether a task is better suited for GPUs, CPUs, or OpenMP runtime. This choice depends on factors such as task size, GPU memory, and kernel launch latency. For example, tiny tasks that require very little computation may prefer to run on CPUs due to the GPU kernel launch overhead, whereas tasks with large memory footprints might prefer to be assigned to OpenMP or CPU when GPU memory is insufficient.
@@
Next, the memory placement of data across different memory spaces is crucial to performance. A mapper must decide where to place data — in the GPU’s FrameBuffer for faster access, in ZeroCopy memory for shared access between CPU and GPU, or in CPU system memory for large data. Each choice introduces a trade-off between memory access speed, memory usage, and transfer overhead.
The memory layout option determines the optimal memory arrangement for data structures. Depending on the memory access patterns and the underlying hardware, selecting between a Struct of Arrays (SOA) or an Array of Structures (AOS) layout, along with constraints on array ordering (Fortran-order or C-order) and memory alignment, can significantly impact performance due to cache efficiency and data locality. Such choice is task-dependent and processor-dependent.
Finally, mappers need to decide how to perform index mapping. As shown in Figure 2, index mapping controls how the data (or more accurately, parallel for loops of task launches) is partitioned and mapped to the processors. The data (or tensors) can be multi-dimensional, and the distributed machines can also be viewed as a processor space (nodes and processors within the node). Index mapping controls the mapping between the two index spaces. Prior work (Unger et al., 2022; Zheng et al., 2022a) shows that searching over how data is partitioned and mapped can significantly change inter-processor communication volume, which affects performance. Commonly-used index mapping functions are shown in Section A.3.
4 A New Approach: Generation and Optimization in a DSL
4.1 Design a Domain-Specific Language
We designed a domain-specific language for mapper generation. We distill the key performance-critical aspects of mapping into language constructs that are easily expressed in the DSL. By providing a higher-level abstraction than C++ APIs, the DSL simplifies interfacing with LLMs. Also, such a modular design enables LLMs to efficiently address the complex optimization challenges of generating high-performance mappers. To implement our DSL, we develop a compiler that can translate the mapper written in DSL into low-level C++ mapping APIs.
Figure 3(a) presents an example DSL mapper that illustrates the key features of our domain-specific mapping language. By contrast, Figure 3(b) shows the (already simplified) code snippet extracted from a C++ mapper. Notably, the C++ snippet covers only part of the functionality that the IndexMap statement (which we discuss in detail below) in the DSL provides.
While a typical C++ mapper requires approximately 400 lines of code, the corresponding DSL mapper can achieve the same functionality in just about 30 lines on average — reflecting a 14×14$\times$14 × reduction, as detailed in Table 1 (based on the 9 applications used in Section 5.2 and Section 5.3). This substantial reduction in code complexity makes the DSL a more accessible target for LLM-based code generation, as it abstracts away many low-level implementation details.
@@ 4.1 Design a Domain-Specific Language
(a) An example mapper in DSL
@@ 9 Rect<2> blocks(zeroes, num_blocks - ones);
10 ...// **126 lines of C++ code ommitted**

### v2

# Improving Parallel Program Performance with LLM Optimizers via Agent-System Interface

@@ Abstract
Modern scientific discovery increasingly relies on high-performance computing for complex modeling and simulation. A key challenge in improving parallel program performance is efficiently mapping tasks to processors and data to memory, a process dictated by intricate, low-level system code known as mappers. Developing high-performance mappers demands days of manual tuning, posing a significant barrier for domain scientists without systems expertise. We introduce a framework that automates mapper development with generative optimization, leveraging richer feedback beyond scalar performance metrics. Our approach features the Agent-System Interface, which includes a Domain-Specific Language (DSL) to abstract away low-level complexity of system code and define a structured search space, as well as AutoGuide, a mechanism that interprets raw execution output into actionable feedback. Unlike traditional reinforcement learning methods such as OpenTuner, which rely solely on scalar feedback, our method finds superior mappers in far fewer iterations. With just 10 iterations, it outperforms OpenTuner even after 1000 iterations, achieving 3.8×3.8$\times$3.8 × faster performance. Our approach finds mappers that surpass expert-written mappers by up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks while reducing tuning time from days to minutes.
Machine Learning, ICML
@@ 1 Introduction
Modern scientific discovery depends on advanced software tools for modeling and simulation (Stocks et al., 2024; Wang et al., 2024; Ltaief et al., 2024). Computational scientists, including physicists, chemists, and biologists, rely on high-performance computing to tackle complex problems. These scientific computations dominate workloads on the world’s most powerful supercomputers (Exascale Computing, ). However, many domain scientists lack expertise in computer science, and therefore having difficulties in optimizing their programs because of the complexity and scale of the underlying machines. Even for experts, finding and fixing performance problems resulting from program modifications or when porting to a new machine is often time-consuming. Any progress on automating performance tuning is of great benefit in this domain.
Task-based programming (Slaughter et al., 2015; Bauer et al., 2012; Augonnet et al., 2009; Chamberlain et al., 2007; Moritz et al., 2018; Barham et al., 2022) has emerged as a promising approach to high performance computing. The paradigm involves decomposing computations into independent tasks that communicate exclusively through their arguments. A key advantage of task-based systems is that the performance tuning problem is factored out into a separate mapping: an assignment of tasks to processors and data to particular memories. High-quality mapping, achieved through a well-designed mapper (implemented as code), can significantly improve performance, often by an order of magnitude (Galvez et al., 2017).
However, currently writing mappers remains a labor-intensive process, as it requires deep knowledge of applications, hardware, and low-level system APIs. In addition, this process is highly application-specific, input-specific, and machine-specific, often taking experts several days of meticulous tuning to achieve high performance. This challenge is especially pronounced for domain scientists, who typically lack the necessary expertise in computer systems and code optimization. Automating mapper development would enable scientists to focus on their own domain of expertise while fully utilizing the capabilities of high-performance computing systems.
Figure 1: Iterative mapper refinement with agent-based generative optimization. The system leverages the Agent-System Interface, which consists of the Domain-Specific Language (DSL) and AutoGuide. The DSL abstracts away the low-level system code, defining a search space for mapping strategies, while AutoGuide interprets execution results into actionable guidance. As iterations progress, the mapper evolves to improve performance.
In this paper, we introduce a system powered by large language models (LLMs) to automate both the generation and optimization of mapper code. The first challenge stems from the complexity of generating mapper code due to the original low-level programming system, which exposes the agent to intricate system APIs, coupled with the problem that raw feedback messages from the system are often uninformative to the agent. The second challenge involves optimizing mapper performance. Specifically, it consists of (1) defining an appropriate search space and (2) devising efficient methods to find optimal mappers, thereby maximizing parallel program performance.
To address the first challenge, we propose an Agent-System Interface (ASI), as shown in Figure 1, an abstraction layer between the agent and the system that simplifies code generation and provides more meaningful feedback to the agent. At the core of ASI is a Domain-Specific Language (DSL), a high-level interface that encapsulates all performance-critical decisions required to generate a mapper. The DSL abstracts away the complexity of low-level system code with a compiler. Additionally, the DSL defines a structured search space, enabling systematic exploration of mapping strategies. We also design and implement the AutoGuide mechanism to interpret raw execution output into informative and actionable guidance. This mechanism allows the agent to iteratively optimize the mapper by leveraging enriched feedback to update its strategy.
For the second challenge, we adopt the generative optimization approach, a recent advance in optimization techniques. Unlike traditional methods such as reinforcement learning (Ansel et al., 2014), which rely solely on scalar rewards, generative optimization can utilize richer forms of feedback, such as error explanations and actionable suggestions expressed in natural language. This agentic optimization workflow has previously proven to be effective across various domains (Nie et al., 2024; Cheng et al., 2024; Yang et al., ; Khattab et al., 2023; Yuksekgonul et al., 2024). Our work is the first to apply such technique to the domain of system optimization.
Our experiments demonstrate that mappers optimized by LLM-powered agents not only match but often surpass expert-written mappers, achieving up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks. Since expert-written mappers set the highest standard, surpassing them is a notable accomplishment. At the same time, our method significantly reduces mapper tuning time from days to minutes, making high-performance mapping more accessible to domain scientists. To further highlight the advantage of generative optimization, we compare it against OpenTuner, a reinforcement learning-based autotuning framework. Our generative optimizer finds mappers $\times1$ faster than OpenTuner when both run for 10 iterations and still maintains a 3.8×3.8$\times$3.8 × advantage even when OpenTuner runs for 1000 iterations. Furthermore, ablation studies underscore the necessity of the agent-system interface design in achieving these performance gains. Our contributions are as follows:
@@
    Design of an Agent-System Interface: We introduce an abstraction layer that simplifies mapper code generation and provides guidance to the agent. The Domain-Specific Language (DSL) defines a search space, allowing the agent to explore mapping strategies without dealing with low-level system code. AutoGuide interprets raw execution output into targeted feedback, enabling the agent to refine mapper code more effectively.
@@
    Generative Optimization for Systems: We introduce generative optimization to improve system performance, leveraging richer feedback such as error messages and actionable suggestions in natural language. Unlike reinforcement learning methods like OpenTuner, which rely solely on scalar feedback, our method identifies better mappers in far fewer iterations. With only 10 iterations, it outperforms OpenTuner by 3.8×3.8$\times$3.8 × even after 1000 iterations.
@@
    Empirical Evaluation of Performance: Our agent-based solution achieves up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks, surpassing expert-written mappers while reducing tuning time from days to minutes. We highlight the critical role of the agent-system interface through ablation studies, demonstrating its impact on achieving the performance gains.
@@ 2 Related Work
Mapping in Parallel Programming Many parallel programming systems allow users to make their own mapping decisions, such as Legion (Bauer et al., 2012), StarPU (Augonnet et al., 2009, 2010), Chapel (Chamberlain et al., 2007), HPX (Kaiser et al., 2014; Heller et al., 2017), Sequoia (Fatahalian et al., 2006), Ray (Moritz et al., 2018), TaskFlow (Huang et al., 2021), and Pathways (Barham et al., 2022). Several techniques have been proposed to automate mapping, including machine learning models (O’Boyle et al., 2013; Wang & O’Boyle, 2009), static analysis (Poesia et al., 2017; Ren et al., 2008), reinforcement learning (Ansel et al., 2014; Mirhoseini et al., 2017) and auto-tuning (SFX Teixeira et al., 2023). We use an agent-based approach with LLMs and explore a larger search space for mappers than traditional methods.
Agentic Frameworks Agents powered by Large Language Models (LLMs) play a critical role in decision-making, planning, tool integration, and solving complex problems in dynamic environments (Guo et al., 2024). Many agentic frameworks have been developed (Yao et al., 2022; Wu et al., 2023; Li et al., 2023; Hong et al., 2023), with uses spanning domains such as software engineering (Gur et al., 2023; Yang et al., 2024b; Jin et al., 2024), robotics (Kannan et al., 2024), healthcare (Li et al., 2024), education (Ramirez & Esparrell, 2024), and knowledge engineering (Shao et al., 2024). Our work is the first to apply an agentic workflow to iteratively optimize mapper code, improving the performance of parallel programs.
AI for Systems The application of AI to optimize system design has gained significant traction in recent years. Techniques such as deep learning (Zheng et al., 2020a, b, 2022b) and gradient-boosted trees (Feng et al., 2023) have been used to predict program execution times for performance optimization. Reinforcement learning methods have addressed challenges in chip floorplanning (Mirhoseini et al., 2021), autotuning (Ansel et al., 2014), auto-vectorization (Haj-Ali et al., 2020a), and compiler phase ordering (Haj-Ali et al., 2020b). While previous efforts have predominantly relied on traditional approaches for cost prediction and optimization, our work uses the recent advances in generative optimization to tackle complex system challenges.
Generative Optimization
Recent work has explored the use of LLMs for optimization problems traditionally tackled with numerical methods, including mixed-integer programming (AhmadiTeshnizi et al., 2024a, b) and numerical optimization (Nie et al., 2024). A key advantage of generative optimization is its ability to iteratively refine solutions using diverse forms of feedback. For example, Cheng et al. (2024) applies generative optimization to robotic manipulation and game playing, while Yuksekgonul et al. (2024) optimizes prompts and molecular designs. While reinforcement learning has been applied to system optimization, the potential of LLM-driven optimization in systems remains unexplored. Our work explores whether generative optimization with richer feedback outperforms traditional methods using scalar rewards in system optimization.
3 Problem Definition
Motivation and Challenges
The concrete problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012). Mappers dictate task scheduling and data placement. A well-designed mapper can achieve orders-of-magnitude speedup over naive strategies.
However, automating mapper generation is challenging due to two key factors. First, the complexity of low-level system code. Implementing a mapper requires writing hundreds of lines of intricate C++ code, demanding expertise in system internals. Second, the vast search space of mapping strategies. The search space grows exponentially with the number of tasks and arguments.
Search Space and Performance Impact
@@ 3 Problem Definition
As illustrated in Figure A1, the search space of mappers involves multiple decisions, each influencing performance. The first key aspect is processor selection, which determines whether a task runs on GPUs, CPUs, or the OpenMP runtime. This choice depends on factors such as task size, GPU memory capacity, and kernel launch overhead. For instance, small tasks may prefer CPUs due to the overhead of launching GPU kernels, while tasks with large memory footprints may run on CPUs when GPU memory is insufficient.
@@
Another crucial dimension is memory placement, which dictates where data is stored. A mapper must decide whether to place data in the GPU’s FrameBuffer for fast access, ZeroCopy memory for CPU-GPU sharing, or CPU system memory for more available storage. Each option presents trade-offs between access speed, memory usage, and data transfer overhead.
Additionally, memory layout further expands the search space, with decisions on Struct of Arrays (SOA) vs. Array of Structures (AOS), data ordering (Fortran-order vs. C-order), and alignment constraints (e.g., 128-byte alignment) significantly affecting cache efficiency and performance.
Finally, an important idiom in high-performance computing is launching tasks over partitioned data. Index mapping determines how data partitions and task executions are distributed across multiple processors. For consistency, we can represent data partitioning as a tensor of data partitions, the machine as a tensor of processors, and tasks operating on the partitioned data as a tensor of tasks. The way data and task indices are mapped to processor indices affects inter-processor communication, a key factor in performance (Unger et al., 2022; Zheng et al., 2022a).
4 Our Approach: Agent-System Interface
@@
(a) An example mapper in Domain-Specific Language (DSL)
@@ 9 Rect<2> blocks(zeroes, num_blocks - ones);
10 ... // 126 lines of C++ code omitted here